OpenStudy (anonymous):

Write an equation of a line that passes through the point (2,3) and is parallel to the line y=3/2x+5 A) y=-2/3x B) y=-2/3x-7 C) y=3/2x D) y=3/2x+7

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

What's the slope of y=3/2x+5

OpenStudy (anonymous):

im not sure

jimthompson5910 (jim_thompson5910):

hint: y = mx+b has a slope of m

OpenStudy (anonymous):

am I supposed to add?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

for example y = 3x + 7 has a slope of 3 notice how m = 3 in this case

OpenStudy (anonymous):

the slop is 3/2?

OpenStudy (anonymous):

* slope

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

the slope of any parallel line to y=3/2x+5 will also have a slope of 3/2 this is because parallel lines have equal slopes

jimthompson5910 (jim_thompson5910):

which choices have slopes of 3/2?

OpenStudy (anonymous):

C & D? @jim_thompson5910

OpenStudy (anonymous):

I think is it C.

jimthompson5910 (jim_thompson5910):

why C and not D?